# Hidden Population III Descendants in Ultra-Faint Dwarf Galaxies

### v1

@@ 3 The energy distribution of Pop III SNe: impact
The models ``````` `only faint”, and ``````` `only hyper” have a high $\chi^{2}>1$, and are thus disfavored. Since faint SNe produce large amounts of C and small of Fe, they result in a long low-Fe tail in the MDF which is not observed in Boötes I (see Fig. 2), and therefore the ``````` `only faint” model is less favoured. On the other hand, assuming that all Pop III SNe explode as hypernovae shifts the peak of the MDF towards lower [Fe/H] values. Indeed, due to the low potential well of Boötes I, when all Pop III explode as energetic hypernovae the majority of the gas and metals are evacuated out of the galaxy. As a result, retaining metals in the halo becomes challenging. In addition, due to the lack of available gas for star formation, the star formation rate of subsequent Pop II is very low ($\Psi\approx 10^{-3}\rm M_{\odot}yr^{-1}$), leading to less efficient metal enrichment.
The other three models give $\chi^{2}<1$. The lowest $\chi^{2}$ value is for the ``````` `only high” model, however, the difference in $\chi^{2}$ is not enought to distinguish between models: $\chi^{2}=0.9$ (only cc), $\chi^{2}=0.4$ (only high), and $\chi^{2}=0.5$ (all types). Therefore, we need to include other observables of the Boötes I UFD to select our “fiducial" model.
$\rm\log(L_{\star}/L_{\odot})$ $\langle\rm[Fe/H]\rangle$ $\chi^{2}$
Boötes I $4.50\pm 0.10$ $-2.58\pm 0.43$
FAINT $4.60\pm 0.01$ $-3.00\pm 0.01$ 1.5
CC $4.59\pm 0.02$ $-2.60\pm 0.01$ 0.9
HIGH $4.60\pm 0.01$ $-2.68\pm 0.03$ 0.4
HYPER $4.40\pm 0.01$ $-3.20\pm 0.01$ 4.8
all TYPES $4.54\pm 0.04$ $-2.70\pm 0.01$ 0.5
Table 1: Observed properties of Boötes I (top row; Kirby et al. 2013; Norris et al. 2010; Lai et al. 2011; Gilmore et al. 2013), and the simulated properties with their corresponding $\chi^{2}$ values, by adopting the different EDFs for Pop III SNe from Fig. 1.
Figure 2: Comparison between the observed (black histograms) and the simulated MDFs of Boötes I. The top four simulated MDFs are obtained assuming that Pop III SNe are all of a single type: faint (orange), core-collape (cc, red), high (light blue), hyper (dark blue). Bottom panel shows the simulated MDF obtained assuming a flat EDF spread among all energies (gray). In each panel the goodness of the fit, $\chi^{2}$, is shown (see also Table 1).
@@
Figure 3: The [C/Fe] (left panel) and [Zn/Fe] (right panel) abundance ratio of Boötes I stars (yellow stars) as a function of their [Fe/H] (SAGA Database Suda et al. (2008)). Colored contours identify the the stellar populations enriched to a level higher than $75\%$ by Pop III SNe with different energies, faint (orange), cc (red), high (lightblue), hyper (dark blue). For comparison we also show the expected chemical abundance ratios of stellar population mainly enriched by normal Pop II SNe (dark purple shaded area). Note that the scales on the y-axis are different in the two panels.
In Fig. 3 we analyze how the different EDFs for Pop III SNe affect the predicted abundance ratios333$\rm[X/Fe]=\log(N_{X}/N_{Fe})-\log(N_{X,\odot}/N_{Fe,\odot})$; where $\rm N_{X}$ and $\rm N_{Fe}$ are the abundances of elements X and Fe, and $\rm N_{X,\odot}$ and $\rm N_{Fe,\odot}$ are the solar values (Asplund et al., 2009) of present-day stars in Boötes I. In particular we focus on [C/Fe] and [Zn/Fe], which are very sensitive to different explosion energies. For each model, the contours delineate the stellar populations that have been predominantly imprinted by a unique type of Pop III SNe, i.e., that contributed $>75\%$ of their present-day mass in metals.
@@
The location of these Pop III descendants in the [X/Fe]--[Fe/H] space depends on both the energy and on the progenitor mass of the Pop III SNe that enriched their birth clouds (see also Vanni et al. 2023). However, we can notice that the descendants of low-energy Pop III SNe (faint, cc) cover a wide range in both [Fe/H] and [C/Fe], $\rm-7.5\lesssim[Fe/H]\lesssim 0$, and, $\rm-1\lesssim[C/Fe]\lesssim 6$, and they are characterized by $\rm[Zn/Fe]\lesssim 0$. On the other hand, Pop III descendants which are predominantly enriched by energetic SNe (high and hypernovae) exhibit a narrower range in $\rm-1\lesssim[C/Fe]\lesssim 3$ and $\rm-4<[Fe/H]<0$. Note that only stars imprinted by high energy SNe and hypernovae can reach super-solar [Zn/Fe] values, up to $\approx+1$ dex. Finally, in Fig.3 we also show the regions covered by stars enriched by normal Pop II SNe, which typically have $\rm[C/Fe]\lesssim 1$ and they can reach high values of $\rm[Zn/Fe]\gtrsim 0$, comparable with those obtained with hypernovae enrichment.
@@
Which EDF for Pop III SNe can reproduce the observed [C/Fe] and [Zn/Fe] ratios of Boötes I stars? From comparing our model results with observational data for Boötes I (Fig. 3) it is evident that none of the individual models which assume a single Pop III SNe type (only faint/cc/high/hyper) is able to reproduce, at the same time, the observed distributions of [C/Fe] and [Zn/Fe]. In fact, assuming that all pristine SNe evolve as low-energy SNe allows to reproduce the high values of [C/Fe] observed in Boötes I stars but not the high [Zn/Fe] values. On the other hand, energetic SNe can reproduce the measured values of [Zn/Fe], but fail to reproduce the observed [C/Fe] stellar ratios. Finally, even taking into account the contribution of Pop II SNe, it is not possible to reproduce the data, as well as assuming a single type of Pop III SNe.
In conclusion, to reproduce the variety of [C/Fe] and [Zn/Fe] measured in Boötes I stars with different [Fe/H] we need the contribution of Pop III SNe with different energies. Thus, we choose as our ``````` `fiducial model” the ``````` `all types”  model (see Fig. 1) in which the theoretical EDF is spread among different equiprobable energies.
@@
Figure 4: Density maps, [X/Fe] --[Fe/H], of the simulated stellar populations at present day, $z=0$. Yellow star symbols are observed chemical abundances of Boötes I stars (SAGA Database Suda et al. (2008)).Top marginal histograms show the MDFs of: observed Boötes I stars (yellow), all stars (light grey), stars enriched at level higher than $75\%$ by Pop III stars (dark purple, top left) and stars enriched at level higher than $75\%$ by Pop II stars (light purple, top right). Right marginal plots show the [X/Fe] abundance ratio distributions for: all simulated stars (light grey), and predominantly ($>75\%$) Pop III/Pop II enriched stars (dark/light purple). The shaded yellow area highlights $\rm[Fe/H]>-4.$
@@ 4 Early chemical evolution: predictions
In Fig. 4, we show all the predicted [X/Fe] versus [Fe/H] density maps for the present-day stellar populations in Boötes I predicted by our fiducial model (``````` `all types” in Fig. 1). The first thing to note is that all the simulated [X/Fe] density maps effectively reproduce the observed data. Indeed, the peaks of the simulated density distributions for each chemical element perfectly match the corresponding observed distributions. It is worth mentioning that the simulated density maps include many more stars than those observed, and cover a broader range in the [X/Fe]-[Fe/H] space. Our model predicts regions of low density ($<20\%$ of the density peak), representing rare stellar populations, not yet observed in UFDs, that are particularly intriguing for investigating the signatures of the first stars.
In the top marginal plots of Fig. 4 we compare the normalized MDFs of Boötes I observed and simulated, including the predicted MDFs for stars predominantly enriched by Pop III ($>75\%$ of their metals, top right) and by Pop II stars (top left). In the range where data are available, $\rm[Fe/H]>-4$, the simulated MDF is dominated by stars predominantly enriched by normal Pop II SNe. Indeed, Pop II enriched stars represent $>90\%$ of the total number of stars in these [Fe/H] range, while stars mainly imprinted by Pop III SNe are extremely rare at $\rm[Fe/H]>-4$, only counting for $\approx 1\%$ of the total population. On the other hand, in the low-metallicity tail of the MDF ($\rm[Fe/H]<-4$), the number of stars that inherited $>75\%$ of their metals from Pop III SNe is of the same order of magnitude as those predominantly enriched by Pop II. As we shall now see, for $\rm[Fe/H]<-4$, Pop III descendants can be distinguished from Pop II enriched stars as outliers, i.e., thanks to their key [X/Fe] abundance ratios.
@@ 4.1 Key abundance ratios to unveil Pop III descendants
Now we investigate if there are regions in the [X/Fe] -- [Fe/H] space where it is possible to uniquely identify the fingerprints of Pop III SNe. To this aim we derive the X-Distribution Function (XDF), i.e. the number of stars in bins of [X/Fe] averaged over different runs, $\langle\rm N_{\star}\rangle$. In the right marginal plots of Fig. 4, we show the comparison between the total XDF (grey), and the XDFs of stars predominantly (>75% of their metals) enriched by Pop III (dark purple) or normal Pop II/I stars (light purple).
For all chemical abundance ratios, the peak of the XDF corresponds to the density peak in star formation and thus it is dominated by stellar populations mainly enriched by normal Pop II/I stars as already explained (see also Rossi et al., 2021). In the low density regions, however, the fingerprints of the first SNe arise, i.e. the XDFs are typically dominated by Pop III enriched stars. In conclusion, in the low metallicity regime, $\rm[Fe/H]<-4$, stars characterized by $\rm[C/Fe]\gtrsim+1$ or $\rm[C/Fe]\lesssim 0$, $\rm[Mg/Fe]\gtrsim+0.5$, $\rm[Na/Fe]\gtrsim+0.5$, $\rm[Si/Fe]\gtrsim+1$, $\rm[Ca/Fe]\gtrsim+0.5$, $\rm[Ni/Fe]\gtrsim+0.5$, or $\rm[Zn/Fe]\lesssim-1$ have a high probability ($>80\%$) to be Pop III descendants.
Figure 5: The [C/Fe] vs [Fe/H] diagram of stellar Boötes I stellar populations mainly enriched by Pop III SNe with different energy: low-energy (first column), high-energy (second column). Top marginal plots show the average number of stars, $\rm\langle N_{\star}\rangle$, as a function of their [Fe/H] of CEMP (solid) and C-normal (dashed) stars enriched at level higher than $75\%$ by different Pop III SNe types, as well as the distribution of overall stars in Boötes I while right marginal plots show $\langle N_{\star}\rangle$ as a function of [C/Fe]. Yellow stars represent the Boötes I data, while double marked points identify stars for which abundance pattern is available in the literature.
@@ 4.2 The imprints of different Pop III SNe types
According to our model predictions, a large fraction of stars with $\rm[C/Fe]\gtrsim+1$ or $\rm[C/Fe]\lesssim 0$ are likely to be Pop III descendants (Fig. 4). Now we go deeper into understanding the origin of these Pop III descendants, exploring the type of SNe that enriched their birth cloud. In Fig. 5 we present the [C/Fe]-[Fe/H] diagram for Boötes I stellar populations predominantly enriched by different Pop III SNe types. As we can see, Pop III descendants can cover a very broad range of both $-7<\rm[Fe/H]<-1$ and $-1\lesssim\rm[C/Fe]\lesssim 4$. However, the probability to find them strongly depends on their location in the [C/Fe]-[Fe/H] space, being maximum for $\rm[Fe/H]\lesssim-3.5$ and $\rm[C/Fe]\gtrsim+1$; or when $\rm[Fe/H]\gtrsim-3.5$ and $\rm[C/Fe]\lesssim 0$. This is also illustrated in Fig. 10 (Appendix A), where we show the absolute carbon abundance, $\rm A(C)$, with respect to $\rm[Fe/H]$.
At $\rm[Fe/H]\lesssim-3.5$, around $90\%$ of the stars in Boötes I are predicted to be CEMP stars, in which the carbon enhancement can be driven by both AGB Pop II stars and Pop III SNe. However, Pop III enriched CEMP stars can be identified through their A(C) (see Rossi et al. (2023) and Appendix A). By inspecting the top marginal plots of Fig.5, where the average number of stars, $\langle\rm N_{\star}\rangle$, predominantly imprinted by Pop III SNe with different energies is shown, it is evident that the enrichment of CEMP stars can arise from different Pop III sources: faint, cc and high energy SNe. Thus, we can conclude that the progenitors of Pop III enriched CEMP stars in Boötes I can be both low-energy (faint, cc) and high-energy Pop III SNe. Conversely, direct descendants of hypernovae are uncommon within this range, $\rm[Fe/H]\lesssim-3.5$, and they mainly have [C/Fe]$<0.7$.
As the [Fe/H] increases, the population of CEMP stars enriched by low-energy Pop III SNe, decreases; while the population of C-normal stars, enriched by energetic (high-energy and hyper) Pop III SNe, increases (see right marginal plot of Fig.5). Note that, towards high [Fe/H] values, the probability to find descendants of Pop III stars declines due to the chemical pollution of subsequent Pop II stellar populations. Therefore, we expect that the majority of stars with $\rm[Fe/H]\gtrsim-3.5$ have originated from gas that has been enriched by multiple generations of Pop II stars (see Appendix A). Nevertheless, among C-normal stars with $\rm[C/Fe]\lesssim+0.7$ we expect to find the imprints of Pop III SNe of different energies, while the imprint of hypernovae is dominant at $\rm[C/Fe]\lesssim-0.5$, accounting for $6\%$ of the total stars in this range.

### v2

@@ 4 Fiducial Model
Different combinations of Pop III EDFs and Pop II/I yields produce MDFs that vary significantly in shape and width. For instance, models assuming “only faint” PopIII SNe produce MDFs spanning a broader range of [Fe/H], shifted toward lower metallicities. Indeed faint SNe produce large amounts of C and small of Fe, resulting in a long low-Fe tail in the MDF which is not observed in Boötes I (first row in Fig. 2), and therefore the ``````` `only faint” model is less favoured. On the other hand, assuming that all Pop III SNe explode as hypernovae shifts the peak of the MDF towards lower [Fe/H] values. Indeed, due to the low potential well of Boötes I, when all Pop III explode as energetic hypernovae the majority of the gas and metals are evacuated out of the galaxy. As a result, retaining metals in the halo becomes challenging. In addition, due to the lack of available gas for star formation, the star formation rate of subsequent Pop II is very low ($\Psi\approx 10^{-3}\rm M_{\odot}yr^{-1}$), leading to less efficient metal enrichment. It is worth noting that the “only high” models, where all Pop III SNe explode as high-energy events, produce MDFs that at first glance appear to match the observed MDF. However, these models perform worse in reproducing other key observables of Boötes I and are therefore strongly disfavored overall.
@@
Figure 3: The [C/Fe] (left) and [Zn/Fe] (right) abundance ratio of observed Boötes I stars (yellow stars) as a function of their [Fe/H] (SAGA Database; Suda et al. 2008). Top row shows the stellar populations enriched by Pop III SNe with different energies: faint (orange), cc (red), high (light blue), hyper (dark blue). Bottom row shows the expected chemical enrichement by normal Pop II/I SNe using different set of stellar yields (LC, NK CC, NK CC+HN). Note that the scales on the y-axis are different in the panels.
The “all types” models, which assume an equal contribution from all Pop III types (faint, core-collapse, high-energy, and hypernovae), provides a more balanced distribution, successfully reproducing the intermediate metallicity range ($-3\lesssim\rm[Fe/H]\lesssim-2$) observed in Boötes I, independently of the Pop II stellar yields adopted.
The choice of Pop II/I stellar yields has an even more pronounced effect, underscoring the dominant role of Pop II/I stars in driving the chemical evolution of Boötes I. For example, the LC yield set systematically produces lower $\chi^{2}$ values compared to NK CC and NK CC+HN, indicating a better agreement with the observed MDF and other constraints. This behavior can be traced back to the higher metal yields — particularly of iron — predicted by the NK yield set compared to LC. In order to reproduce the shape and normalization of the observed MDF, which carries the highest statistical weight in the total $\chi^{2}$ evaluation, the models compensate by adjusting the star formation efficiency ($\epsilon_{\star}$) and the outflow efficiency ($\epsilon_{w}$). Specifically, the enhanced Fe production per SN event in NK models generally requires a lower $\epsilon_{\star}$ to avoid an overproduction of metals. However, this leads to reduced stellar mass growth and lower predicted star counts, often resulting in tension with other constraints, such as the observed stellar luminosity. The overall worse performance of the NK-based models is therefore a consequence of the interplay between their nucleosynthetic yields and the global galactic parameters required to reproduce the MDF. Among the tested combinations, the “all types” Pop III EDF paired with the LC Pop II/I yields emerges as the best-fit model, achieving the lowest reduced $\chi^{2}$ value and providing a realistic reproduction of the observed properties of Boötes I.
@@
In Fig. 3 we analyze how the different EDFs for Pop III SNe affect the predicted abundance ratios444$\rm[X/Fe]=\log(N_{X}/N_{Fe})-\log(N_{X,\odot}/N_{Fe,\odot})$; where $\rm N_{X}$ and $\rm N_{Fe}$ are the abundances of elements X and Fe, and $\rm N_{X,\odot}$ and $\rm N_{Fe,\odot}$ are the solar values (Asplund et al., 2009) of present-day stars in Boötes I. In particular we focus on [C/Fe] and [Zn/Fe], which are very sensitive to different explosion energies. For each model, the contours delineate the stellar populations that have been predominantly imprinted by a unique type of Pop III SNe, i.e., that contributed $>95\%$ of their present-day mass in metals. The location of these Pop III descendants in the [X/Fe]--[Fe/H] space depends on both the energy and on the progenitor mass of the Pop III SNe that enriched their birth clouds (see also Vanni et al. 2023). However, we can notice that the descendants of low-energy Pop III SNe (faint, cc) cover a wide range in both [Fe/H] and [C/Fe], $\rm-7.5\lesssim[Fe/H]\lesssim-1$, and, $\rm-1\lesssim[C/Fe]\lesssim 6$, and they are characterized by $\rm[Zn/Fe]\lesssim 0$. On the other hand, Pop III descendants which are predominantly enriched by energetic SNe (high and hypernovae) exhibit a narrower range in $\rm-1\lesssim[C/Fe]\lesssim 3$ and $\rm-4<[Fe/H]<0$. Note that only stars imprinted by high energy SNe and hypernovae can reach super-solar [Zn/Fe] values, up to $\approx+1$ dex. Finally, in Fig.3 (bottom row) we also show the regions covered by stars enriched by normal Pop II/I SNe assuming different set of stellar yields, which have $\rm[C/Fe]\lesssim 1$ and they can reach high values of $\rm[Zn/Fe]\gtrsim 0$, comparable with those obtained with hypernovae enrichment.
@@
From comparing our model results with observational data for Boötes I (Fig. 3) it is evident that none of the individual models which assume a single Pop III SNe type (only faint/cc/high/hyper) is able to reproduce, at the same time, the observed distributions of [C/Fe] and [Zn/Fe]. In fact, assuming that all pristine SNe evolve as low-energy SNe allows to reproduce the high values of [C/Fe] observed in Boötes I stars but not the high [Zn/Fe] values. On the other hand, energetic SNe can reproduce the measured values of [Zn/Fe], but fail to reproduce the observed [C/Fe] stellar ratios. Finally, even taking into account the contribution of Pop II SNe and using different set of stellar yields/EDFs, it is not possible to reproduce the data, as well as assuming a single type of Pop III SNe. In conclusion, to reproduce the variety of [C/Fe] and [Zn/Fe] measured in Boötes I stars with different [Fe/H] we need the contribution of Pop III SNe with different energies.
Thus, we adopt as our ``````` `fiducial model” the one that minimizes the reduced chi-squared: the ``````` `all types” EDF (see Fig.1), paired with the LC Pop II stellar yields.
@@
Figure 4: Density maps, [X/Fe] --[Fe/H], of the simulated stellar populations at present day, $z=0$. Yellow star symbols are observed chemical abundances of Boötes I stars (SAGA Database Suda et al. (2008)) , white symbols indicate stars for which [X/Fe] represents an upper limit, while red symbols denote potential candidates for Pop III descendants.. Top marginal histograms show the MDFs of:observed Boötes I stars (yellow), all stars (light grey), stars enriched at level higher than $75\%$ by Pop III stars (dark purple, top left) and stars enriched at level higher than $75\%$ by Pop II stars (light purple, top right). Right marginal plots show the [X/Fe] abundance ratio distributions for: all simulated stars (light grey), and predominantly ($>75\%$) Pop III/Pop II enriched stars (dark/light purple). Shaded yellow area highlights $\rm[Fe/H]>-4$.
@@ 5 Early chemical evolution: predictions
In Fig. 4, we show the predicted [X/Fe] versus [Fe/H] density maps for the present-day stellar populations in Boötes I predicted by our fiducial model (``````` `all types” in Fig. 1 top panel).
The first thing to note is that all the simulated [X/Fe] density maps effectively reproduce the observed data, with the exception of [Sc/Fe], where there is a discrepancy of approximately one dex between the observations and simulations. This discrepancy in [Sc/Fe] is a known issue in nucleosynthesis modeling and has been highlighted in previous studies (e.g., Kobayashi et al. (2020)). Indeed, in all other cases the peaks of the simulated density distributions for each chemical element perfectly match the corresponding observed distributions. It is worth mentioning that the simulated density maps include many more stars than those observed, and cover a broader range in the [X/Fe]-[Fe/H] space. Our model predicts regions of low density ($<20\%$ of the density peak), representing rare stellar populations, not yet observed in UFDs, that are particularly intriguing for investigating the signatures of the first stars.
In the top marginal plots of Fig. 4 we compare the normalized MDFs of Boötes I observed and simulated, including the predicted MDFs for stars predominantly enriched by Pop III ($>75\%$ of their metals, top right) and by Pop II stars (top left). In the range where data are available, $\rm[Fe/H]>-4$, the simulated MDF is dominated by stars predominantly enriched by normal Pop II SNe. Indeed, Pop II enriched stars represent $>90\%$ of the total number of stars in these [Fe/H] range, while stars mainly imprinted by Pop III SNe are extremely rare at $\rm[Fe/H]>-4$, only counting for $\approx 1\%$ of the total population. In the low-metallicity tail of the MDF ($\rm[Fe/H]<-4$), more than $80\%$ of the stars are predominantly imprinted by Pop III SNe. These Pop III descendants can be further distinguished from Pop II-enriched stars in this regime by their unique [X/Fe] abundance ratios, marking them as outliers in the chemical abundance space.
@@ 5.1 Key abundance ratios to unveil Pop III descendants
Now we investigate if there are regions in the [X/Fe] -- [Fe/H] space where it is possible to uniquely identify the fingerprints of Pop III SNe. To this aim we derive the X-Distribution Function (XDF), i.e. the number of stars in bins of [X/Fe] averaged over different runs, $\langle\rm N_{\star}\rangle$.
In the right marginal plots of Fig. 4, we show the comparison between the total XDF (grey), and the XDFs of stars predominantly (>75% of their metals) enriched by Pop III (dark purple) or normal Pop II/I stars (light purple).
For all chemical abundance ratios, the peak of the XDF corresponds to the density peak in star formation and thus it is dominated by stellar populations mainly enriched by normal Pop II/I stars as already explained (see also Rossi et al., 2021). In the low density regions, however, the fingerprints of the first SNe arise, i.e. the XDFs are typically dominated by Pop III enriched stars. In conclusion, stars characterized by $\rm[C/Fe]\gtrsim+1$ or $\rm[C/Fe]\lesssim 0$, $\rm[Mg/Fe]\gtrsim+0.5$, $\rm[Na/Fe]\gtrsim+0.5$, $\rm[Si/Fe]\gtrsim+1$, $\rm[Ca/Fe]\gtrsim+0.5$, $\rm[Ni/Fe]\gtrsim+0.5$, or $\rm[Zn/Fe]\lesssim-1$ have a high probability ($>80\%$) to be Pop III descendants.
Figure 5: The [C/Fe] vs [Fe/H] diagram of predicted Boötes I stellar populations mainly enriched (>75%) by Pop III SNe with different energy: low-energy (first column), high-energy (second column). Top marginal plots show the average number of stars, $\rm\langle N_{\star}\rangle$, as a function of [Fe/H] for CEMP (solid) and C-normal (dashed) stars enriched at level higher than $75\%$ by different Pop III SNe types, as well as the overall MDF in Boötes I. Yellow stars represent the Boötes I data, while double marked points identify stars for which abundance pattern is available in the literature. Red stars represent candidates for Pop III descendants.
@@ 5.2 The imprints of different Pop III SNe types
According to our model predictions, a large fraction of stars with $\rm[C/Fe]\gtrsim+1$ (at $\rm[Fe/H]\lesssim-3.5)$ or $\rm[C/Fe]\lesssim 0$ (at $\rm[Fe/H]\gtrsim-3$) are likely to be Pop III descendants (Fig. 4). Now we go deeper into understanding the origin of these Pop III descendants, exploring the type of SNe that enriched their birth cloud. In Fig. 5 we present the [C/Fe]-[Fe/H] diagram for Boötes I stellar populations predominantly enriched by different Pop III SNe types. As we can see, Pop III descendants can cover a very broad range of both $-7<\rm[Fe/H]<-1$ and $-1\lesssim\rm[C/Fe]\lesssim 4$. However, the probability to find them strongly depends on their location in the [C/Fe]-[Fe/H] space, being maximum for $\rm[Fe/H]\lesssim-3.5$ and $\rm[C/Fe]\gtrsim+1$; or when $\rm[Fe/H]\gtrsim-3.5$ and $\rm[C/Fe]\lesssim 0$. This is also illustrated in Fig. 10 (Appendix A), where we show the absolute carbon abundance, $\rm A(C)$, with respect to $\rm[Fe/H]$.
At $\rm[Fe/H]\lesssim-3.5$, around $90\%$ of the stars in Boötes I are predicted to be CEMP stars, in which the carbon enhancement is driven by Pop III SNe. By inspecting the top marginal plots of Fig.5, where the average number of stars, $\langle\rm N_{\star}\rangle$, predominantly imprinted by Pop III SNe with different energies is shown, it is evident that the enrichment of CEMP stars can arise from different Pop III sources: faint, cc and high energy SNe. Thus, we can conclude that the progenitors of Pop III enriched CEMP stars in Boötes I can be both low-energy (faint, cc) and high-energy Pop III SNe.
Conversely, direct descendants of hypernovae are uncommon within this range, $\rm[Fe/H]\lesssim-3.5$, and they mainly have [C/Fe]$<0.7$.
As the [Fe/H] increases, the population of CEMP stars enriched by low-energy Pop III SNe, decreases; while the population of C-normal stars, enriched by energetic (high-energy and hyper) Pop III SNe, increases (see right marginal plot of Fig.5). Note that, towards high [Fe/H] values, the probability to find descendants of Pop III stars declines due to the chemical pollution of subsequent Pop II stellar populations. Therefore, we expect that the majority of stars with $\rm[Fe/H]\gtrsim-3.5$ have originated from gas that has been enriched by multiple generations of Pop II stars (see Appendix A).
Nevertheless, among C-normal stars with $\rm[C/Fe]\lesssim+0.7$ we expect to find the imprints of Pop III SNe of different energies, while the imprint of hypernovae is dominant at $\rm[C/Fe]\lesssim-0.5$, accounting for $6\%$ of the total stars in this range.